Last year the employees of the city health department donated an average of $ 8 to the rescue squad. Test the hypothesis at the 0.01 level of significance that the average contribution this year is still $ 8 if a random sample of 35 employees showed an average donation of $ 8.90 with a standard deviation of $ 1.75.

1
Expert's answer
2022-03-03T09:59:42-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu=8"

"H_1:\mu\not=8"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=34" degrees of freedom, and the critical value for a two-tailed test is "t_c=2.728394."

The rejection region for this two-tailed test is "R = \{t: |t| > 2.728394\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{8.90-8}{1.75/\sqrt{35}}\approx3.042555"

Since it is observed that "|t| = 3.042555 > 2.728394=t_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=34" degrees of freedom, "t=3.042555," is "p=0.0045," and since "p=0.0045<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 8, at the "\alpha = 0.01" significance level.